# Dark Photon Dark Matter and Low-Frequency Gravitational Wave Detection with Gaia-like Astrometry

### v1

@@ V Conclusion
Astrometric data can also be correlated with the current existing PTA data, allowing for a more concrete validation of the correlation pattern induced by GWs or dark matter and a more sensitive probe of macroscopic polarizations of the GWs Qin et al. (2019, 2021); Çalışkan et al. (2024). Compared to PTAs, future astrometric surveys like Roman Wang et al. (2021, 2022) is able to probe $\mu$Hz GWs or beyond and ultralight bosons with higher masses, allowing for cross-correlation with other proposed measurements in the same frequency range such as binary resonance Blas and Jenkins (2022a, b) and Doppler tracking Zwick et al. (2024), potentially closing the gap between PTA and LISA frequency ranges.
@@ Acknowledgement
We would like to thank Huai-Ke Guo for his collaboration in the early stages of this project. We also thank Christopher J. Moore for useful email exchanges. H.A. is supported in part by the National Key R&D Program of China under Grant No. 2023YFA1607104 and 2021YFC2203100, the National Science Foundation of China (NSFC) under Grant No. 11975134, and the Tsinghua University Dushi Program. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No. 12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY7003. X.W. acknowledges the Royal Society as the funding source of the Newton International Fellowship. X.X. is supported by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC2121 “Quantum Universe” — 390833306. Y.Z. is supported by the U.S. Department of Energy under Award No. DESC0009959.
Note added: During the preparation of this paper, we became aware of the recent work by Kim (2024) and Dror and Verner (2024) which considered the astrometric effect of the ultralight scalar dark matter. Our work focuses on the detection of the ultralight vector dark matter and continuous gravitational waves, which is distinctive from their themes.
@@ Appendix A Metric perturbation induced by DPDM
Consider a general metric perturbation $h_{\mu\nu}$, the equation of motion can be written as Carroll (2019)
$\Box h_{\mu\nu}-\partial^{\alpha}\partial_{\nu}h_{\mu\alpha}-\partial^{\alpha}\partial_{\mu}h_{\nu\alpha}+\partial_{\mu}\partial_{\nu}h+\eta_{\mu\nu}(\partial^{\alpha}\partial^{\beta}h_{\alpha\beta}-\Box h)=-16\pi GT_{\mu\nu}\;,$ (23)
@@
$-16\pi GT_{00}$ $=\partial_{0}^{2}h-(\partial_{k}\partial_{l}h_{kl}-\Box h)\;,$ (24)
$-16\pi GT_{0i}$ $=-\partial_{k}\partial_{0}h_{ik}+\partial_{i}\partial_{0}h\;,$ (25)
$-16\pi GT_{ij}$ $=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\delta_{ij}\Box h\;.$ (26)
From Eq. (26) we find
$-2\Box h+\nabla^{2}h+\partial_{l}\partial_{j}h_{jl}=-16\pi GT_{kk}\;.$ (27)
@@
$-16\pi G\left(T_{ij}-\frac{\delta_{ij}}{2}T_{kk}\right)=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\frac{1}{2}\delta_{ij}\nabla^{2}h\;.$ (28)
Next we calculate the stress-energy tensor induced by the DPDM. The virial velocity of our galaxy is $v_{\rm vir}\simeq{\cal O}(10^{-3})$, so for ultralight DPDM with mass $m_{A}\sim\mathrm{10^{-23}~{}eV}$, the de Broglie wavelength is $\lambda_{dB}\sim 4\,\mathrm{kpc}\,(10^{-23}\,\mathrm{eV}/m_{A})(10^{-3}/v_{\rm vir})$, which is of the same order as the galactic size. The coherence time is $t_{c}=4\pi/(m_{A}v^{2}_{\rm vir})=26\,\mathrm{Myr}\,(10^{-23}\,\mathrm{eV}/{m_{A}})(10^{-3}/v_{\rm vir})^{2}$. Since we are considering DPDM within a coherent time, we can write down the spatial component of dark photon field as
$\bm{A}(t,\boldsymbol{x})=A_{0}(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})e^{iEt}\right]\;,$ (29)
@@
$T_{\mu\nu}(t,\boldsymbol{x})=\eta_{\mu\nu}\left(-\frac{1}{4}F_{\rho\sigma}F^{\rho\sigma}-\frac{1}{2}m^{2}A^{\rho}A_{\rho}\right)+\eta^{\rho\sigma}F_{\mu\sigma}F_{\nu\rho}+m^{2}A_{\mu}A_{\nu}\;.$ (31)
@@
$T_{00}^{(0)}=\rho(\boldsymbol{x})\;,\qquad T_{0k}^{(0)}=0\;,$ (32)
$T^{(0)}_{kk}=-\rho(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\;,$
$T_{ij}^{(0)}-\frac{\delta_{ij}}{2}T_{kk}=2\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]e^{2iEt}\right\}\;.$
where for simplicity we defined $\rho(\boldsymbol{x})\equiv m^{2}_{A}A_{0}^{2}(\boldsymbol{x})/2$. By combining Eq. (24) and Eq. (27), we can eliminate $\partial_{l}\partial_{k}h_{kl}$ and get
$\partial_{0}^{2}h=-8\pi G\left(T_{00}+T_{kk}\right)=-8\pi G\rho(\boldsymbol{x})\left\{1-{\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\right\}\;.$ (33)
Integrating Eq. (33), we find that the trace $h$ contains at least an oscillating component
$h^{\rm osc}=-\frac{8\pi G\rho(\boldsymbol{x})}{4m^{2}_{A}}{\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2iEt}\right]\;,$ (34)
which suggests that $h_{ij}$ also has an oscillating part
$h_{ij}^{\rm osc}(t,\boldsymbol{x})=-h_{0}^{\rm osc}(\boldsymbol{x}){\rm Re}\left[\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2iEt}\right]\;.$ (35)
Combining Eq. (28) and Eq. (32), we obtain
$-32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2iEt}\right\}=\Box h_{ij}$ $-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h$
$+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\frac{1}{2}\delta_{ij}\nabla^{2}h\;,$ (36)
which is accurate to the zeroth order of $v_{\rm vir}$. We extract the oscillation parts on both sides of Eq. (36), which yields
$\partial_{0}^{2}{\rm Re}\left[-h_{0}^{\rm osc}(\boldsymbol{x})\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2iEt}\right]=32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2iEt}\right\}\;,$ (37)
@@
$\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})=q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\;.$ (38)
@@
$h_{0}^{\rm osc}(\boldsymbol{x})=\frac{8\pi G\rho(\boldsymbol{x})}{m^{2}_{A}}\;.$ (39)
@@ Appendix B Gauge invariance of the angular deflection
In this appendix we calculate the angular deflection of light from stars due to the oscillating metric perturbations. This part follows the discussion in Book and Flanagan (2011). In this part, we also prove the gauge invariance of the angular deflection. Consider a general metric perturbation
$\mathrm{d}s^{2}=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}=(\eta_{\mu\nu}+h_{\mu\nu})\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}\;.$ (40)
where we work on Minkowski background with $\eta_{\mu\nu}={\rm diag}\{-1,1,1,1\}$. We consider a photon travels from the source to the detector, then the angular deflection of light we need to calculate becomes angular difference of the observed photon momentum compared to the flat spacetime case. Without metric perturbation, we choose the detector to be at the origin $(0,\boldsymbol{0})$ and the unperturbed photon worldline to be $x_{(0)}^{\mu}=\omega_{0}(\lambda,-\lambda\boldsymbol{n})$, where $\boldsymbol{n}$ is the unit vector of the source, $\omega_{0}$ is the unperturbed photon frequency. The unperturbed photon four-momentum is $k_{(0)}^{\mu}=\omega_{0}\mathcal{N}^{\mu}$ where we define $\mathcal{N}^{\mu}=(1,-\boldsymbol{n})$. Suppose the unperturbed position of the source is $x_{s}^{i}(t)=x_{s}^{i}$, then the affine parameter of the source is $\lambda_{s}=-|\boldsymbol{x}_{s}|/\omega_{0}$.
@@
$k^{\mu}(\lambda)=k_{(0)}^{\mu}(\lambda)+k_{(1)}^{\mu}(\lambda)+\mathcal{O}(h^{2})\;,$ (41)
@@
$\frac{\mathrm{d}k^{\rho}}{\mathrm{d}\lambda}=-\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (42)
We integrate from $\lambda=0$ to arbitrary $\lambda$ to get $k^{\mu}(\lambda)$
$k^{\mu}(\lambda)=k_{(0)}^{\mu}+\kappa^{\mu}-\int_{0}^{\lambda}\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime}\;,$ (43)
where $\kappa^{\mu}$ represents the first order perturbation term of the photon four-momentum at $\lambda=0$. Integrating Eq. (43) again gives the full trajectory of photon
$x^{\mu}(\lambda)=\lambda(k_{(0)}^{\mu}+\kappa^{\mu})-\int_{0}^{\lambda}\int_{0}^{\lambda^{\prime}}\Gamma^{\rho}_{\mu\nu}(\lambda^{\prime\prime})k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+x^{\mu}(0)\;,$ (44)
where we use $x^{\mu}(0)$ to express the final position of the photon, which also refers to the position of the detector. Choosing $\lambda=\lambda_{s}$ in Eq. (44), we obtain
$\kappa^{\mu}=\frac{1}{\lambda_{s}}\left(x^{\mu}(\lambda_{s})-x^{\mu}(0)+\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\Gamma^{\mu}_{\gamma\nu}(\lambda^{\prime\prime})k^{\gamma}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\right)-k_{(0)}^{\mu}\;.$ (45)
To the zeroth order of perturbation, $x^{\mu}(\lambda_{s})-x^{\mu}(0)=\lambda_{s}k_{(0)}^{\mu}$, so they should cancel with each other in Eq. (45). However, in a general case with non-zero metric perturbation, we can not assume that the detector and source are static since the geodesic equations of their spatial trajectories also contain $\mathcal{O}(h)$ perturbations. Therefore we keep $x^{\mu}(\lambda_{s})$ and $x^{\mu}(0)$ in the expression of $\kappa^{\mu}$.
In addition, the null geodesic equation constraint $g_{\mu\nu}(\lambda)k^{\mu}(\lambda)k^{\nu}(\lambda)=0$ yields
$\kappa_{0}=n_{i}\kappa_{i}-\frac{1}{2\omega_{0}}h_{\mu\nu}(0)k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (46)
@@
$e^{\nu}_{\alpha}=\delta^{\nu}_{\alpha}+\delta e^{\nu}_{\alpha}\;,\quad\alpha=0,1,2,3\;.$ (47)
@@
$e^{\mu}_{0}\nabla_{\mu}e^{\nu}_{\alpha}=0\;.$ (48)
Inserting Eq. (47) into Eq. (48), we find
$\partial_{0}\delta e^{\nu}_{\alpha}+\Gamma_{\alpha 0}^{\nu}=0\;.$ (49)
@@
$g_{\mu\nu}e^{\nu}_{\alpha}e^{\mu}_{\beta}=\delta_{\alpha\beta}\;.$ (50)
We keep only the first order terms of Eq. (50) which gives another constraint on $\delta e^{\nu}_{\alpha}$
$h_{\mu\nu}+\eta_{\mu\gamma}\delta e^{\gamma}_{\nu}+\eta_{\nu\gamma}\delta e^{\nu}_{\mu}=0\;.$ (51)
@@
$\delta n_{i}=-\frac{g_{\nu\mu}(0)k^{\nu}(0)e^{\mu}_{i}}{\omega_{d}}-n_{i}\;,$ (52)
where $\omega_{d}=-g_{\nu\mu}(0)k^{\nu}$ is the observed photon frequency. Expanding Eq. (52) to the first order, we obtain
$\delta n_{i}=-\frac{1}{\omega_{0}}\left(h_{\nu i}(0)k^{\nu}_{(0)}+\kappa_{i}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{i}-n_{i}(h_{\nu 0}(0)k^{\nu}_{(0)}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{0}+\kappa_{0})\right)\;.$ (53)
Plugging Eq. (46) into Eq. (53), we find
$\delta n_{i}=-\frac{1}{\omega_{0}}P_{ij}\kappa_{j}-h_{\nu i}(0)\mathcal{N}^{\nu}-\mathcal{N}_{\mu}\delta e^{\mu}_{i}+n_{i}\left(h_{\nu 0}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}k_{d(1)}^{\mu}\right)-\frac{1}{2}n_{i}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}\;,$ (54)
where $P_{ij}=\delta_{ij}-n_{i}n_{j}$ is the projection operator. Multiplying both sides of Eq. (54) by $n^{i}$, we obtain
@@
$=0\;.$ (55)
Note that in the last step we use Eq. (51). Now $n^{i}\delta n_{i}=0$ shows that $\delta n_{i}$ is orthogonal to $n^{i}$, so we can multiply $P_{ij}$ to each term in the expression of $\delta n_{i}$, which reads
$\delta n_{i}=-P_{ij}\left(\frac{1}{\omega_{0}}\kappa_{j}+h_{\nu j}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\delta e^{\mu}_{j}\right)\;.$ (56)

### v2

@@ 5 Conclusion
Astrometric data can also be correlated with the current existing PTA data, allowing for a more concrete validation of the correlation pattern induced by GWs or dark matter and a more sensitive probe of macroscopic polarizations of the GWs (Qin et al. (2019, 2021); Çalışkan et al. (2024)). Compared to PTAs, future astrometric surveys like Roman (Wang et al. (2021, 2022)) is able to probe $\mu$Hz GWs or beyond and ultralight bosons with higher masses, allowing for cross-correlation with other proposed measurements in the same frequency range such as binary resonance (Blas & Jenkins (2022a, b)) and Doppler tracking (Zwick et al. (2024)), potentially closing the gap between PTA and LISA frequency ranges.
@@ Acknowledgement
We would like to thank Huai-Ke Guo for his collaboration in the early stages of this project. We also thank Christopher J. Moore for useful email exchanges. H.A. is supported in part by the National Key R&D Program of China under Grant No. 2023YFA1607104 and 2021YFC2203100, the National Natural Science Foundation of China (NSFC) under Grant No. 11975134, and the Tsinghua University Dushi Program. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China under Grants No. 12025507, No. 12150015; and is supported by the Key Research Program of Frontier Science of the Chinese Academy of Sciences (CAS) under Grants No. ZDBS-LY7003. X.W. acknowledges the Royal Society as the funding source of the Newton International Fellowship. X.X. is supported by Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC2121 “Quantum Universe” — 390833306. Y.Z. is supported by the U.S. Department of Energy under Award No. DESC0009959.
Note added: During the preparation of this paper, we became aware of the recent work by (Kim (2024)) and (Dror & Verner (2024)) which considered the astrometric effect of the ultralight scalar dark matter. Our work focuses on the detection of the ultralight vector dark matter and continuous gravitational waves, which is distinctive from their themes.
[…]
@@ Appendix A Metric perturbation induced by DPDM
Consider a general metric perturbation $h_{\mu\nu}$, the equation of motion can be written as (Carroll (2019))
$\Box h_{\mu\nu}-\partial^{\alpha}\partial_{\nu}h_{\mu\alpha}-\partial^{\alpha}\partial_{\mu}h_{\nu\alpha}+\partial_{\mu}\partial_{\nu}h+\eta_{\mu\nu}(\partial^{\alpha}\partial^{\beta}h_{\alpha\beta}-\Box h)=-16\pi GT_{\mu\nu}\;,$ (A1)
@@
$-16\pi GT_{00}$ $=\partial_{0}^{2}h-(\partial_{k}\partial_{l}h_{kl}-\Box h)\;,$ (A2)
$-16\pi GT_{0i}$ $=-\partial_{k}\partial_{0}h_{ik}+\partial_{i}\partial_{0}h\;,$ (A3)
$-16\pi GT_{ij}$ $=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\delta_{ij}\Box h\;.$ (A4)
From Eq. (A4) we find
$-2\Box h+\nabla^{2}h+\partial_{l}\partial_{j}h_{jl}=-16\pi GT_{kk}\;.$ (A5)
@@
$-16\pi G\left(T_{ij}-\frac{\delta_{ij}}{2}T_{kk}\right)=\Box h_{ij}-\partial_{k}\partial_{i}h_{jk}-\partial_{k}\partial_{j}h_{ik}+\partial_{i}\partial_{j}h+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}-\frac{1}{2}\delta_{ij}\nabla^{2}h\;.$ (A6)
Next we calculate the stress-energy tensor induced by the DPDM. The virial velocity of our galaxy is $v_{\rm vir}\simeq{\cal O}(10^{-3})$, so for ultralight DPDM with mass $m_{A}\sim\mathrm{10^{-23}~{}eV}$, the de Broglie wavelength is $\lambda_{dB}\sim 4\,\mathrm{kpc}\,(10^{-23}\,\mathrm{eV}/m_{A})(10^{-3}/v_{\rm vir})$, which is of the same order as the galactic size. The coherence time is $t_{c}\sim 1/(m_{A}v^{2}_{\rm vir})=2\,\mathrm{Myr}\,(10^{-23}\,\mathrm{eV}/{m_{A}})(10^{-3}/v_{\rm vir})^{2}$. Since we are considering DPDM within a coherent time, we can write down the spatial component of dark photon field as
$\bm{A}(t,\boldsymbol{x})=A_{0}(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})e^{iEt}\right]\;,$ (A7)
@@
$T_{\mu\nu}(t,\boldsymbol{x})=\eta_{\mu\nu}\left(-\frac{1}{4}F_{\rho\sigma}F^{\rho\sigma}-\frac{1}{2}m^{2}A^{\rho}A_{\rho}\right)+\eta^{\rho\sigma}F_{\mu\sigma}F_{\nu\rho}+m^{2}A_{\mu}A_{\nu}\;.$ (A9)
@@
$T_{00}^{(0)}=\rho(\boldsymbol{x})\;,\qquad T_{0k}^{(0)}=0\;,$ (A10)
$T^{(0)}_{kk}=-\rho(\boldsymbol{x}){\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2im_{A}t}\right]\;,$
$T_{ij}^{(0)}-\frac{\delta_{ij}}{2}T_{kk}=2\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]e^{2im_{A}t}\right\}\;.$
where for simplicity we defined $\rho(\boldsymbol{x})\equiv m^{2}_{A}A_{0}^{2}(\boldsymbol{x})/2$. By combining Eq. (A2) and Eq. (A5), we can eliminate $\partial_{l}\partial_{k}h_{kl}$ and get (hereafter we add superscript “DP” for quantities associated with the dark photon dark matter)
$\partial_{0}^{2}h^{\rm DP}=-8\pi G\left(T_{00}+T_{kk}\right)=-8\pi G\rho(\boldsymbol{x})\left\{1-{\rm Re}\left[\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\,e^{2im_{A}t}\right]\right\}\;.$ (A11)
Integrating Eq. (A11), we find that the trace $h$ contains at least an oscillating component
$h^{\rm DP}(\boldsymbol{x})=-\frac{8\pi G\rho(\boldsymbol{x})}{m^{2}_{A}}{\rm Re}\left[\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})}{4}\,e^{2im_{A}t}\right]\;,$ (A12)
which suggests that $h_{ij}$ also has an oscillating part and a static part
$h_{ij}^{\rm DP}(t,\boldsymbol{x})=-h_{0}^{\rm DP}(\boldsymbol{x}){\rm Re}\left[\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2im_{A}t}\right]\;.$ (A13)
Combining Eq. (A6) and Eq. (A10), we obtain
$-32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2im_{A}t}\right\}=\Box h_{ij}^{\rm DP}$ $-\partial_{k}\partial_{i}h_{jk}^{\rm DP}-\partial_{k}\partial_{j}h_{ik}^{\rm DP}+\partial_{i}\partial_{j}h^{\rm DP}$
$+\frac{1}{2}\delta_{ij}\partial_{k}\partial_{l}h_{kl}^{\rm DP}-\frac{1}{2}\delta_{ij}\nabla^{2}h^{\rm DP}\;,$ (A14)
which is accurate to the zeroth order of $v_{\rm vir}$. We extract the oscillation parts on both sides of Eq. (A14), which yields
$\partial_{0}^{2}{\rm Re}\left[-h_{0}^{\rm DP}(\boldsymbol{x})\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})\,e^{2im_{A}t}\right]=32\pi G\rho(\boldsymbol{x}){\rm Re}\left\{\left[q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\right]\,e^{2im_{A}t}\right\}\;,$ (A15)
@@
$\mathcal{H}^{\rm DP}_{ij}(\boldsymbol{x})=q_{i}(\boldsymbol{x})q_{j}(\boldsymbol{x})-\frac{\boldsymbol{q}(\boldsymbol{x})\cdot\boldsymbol{q}(\boldsymbol{x})\delta_{ij}}{4}\;.$ (A16)
@@
$h_{0}^{\rm DP}(\boldsymbol{x})=\frac{8\pi G\rho(\boldsymbol{x})}{m^{2}_{A}}\;.$ (A17)
@@ Appendix B Gauge invariance of the angular deflection
In this appendix we calculate the angular deflection of light from stars due to the oscillating metric perturbations. This part follows the discussion in Book & Flanagan (2011). In this part, we also prove the gauge invariance of the angular deflection. Consider a general metric perturbation
$\mathrm{d}s^{2}=g_{\mu\nu}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}=(\eta_{\mu\nu}+h_{\mu\nu})\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}\;.$ (B1)
where we work in the Minkowski background with $\eta_{\mu\nu}={\rm diag}\{-1,1,1,1\}$. Now consider a photon travels from the source to the detector, then the angular deflection of light we need to calculate becomes angular difference of the observed photon momentum compared to the flat spacetime case. Without metric perturbation, we choose the detector to be at the origin $(0,\boldsymbol{0})$ and the unperturbed photon worldline to be $x_{(0)}^{\mu}=\omega_{0}\mathcal{N}^{\mu}\lambda$, where $\boldsymbol{n}$ is the unit vector of the source, $\omega_{0}$ is the unperturbed photon frequency, and $\mathcal{N}^{\mu}=(1,-\boldsymbol{n})$. The unperturbed photon four-momentum is $k_{(0)}^{\mu}=\omega_{0}\mathcal{N}^{\mu}$. Suppose the unperturbed position of the source is $x_{s}^{i}(t)=x_{s}^{i}$, then the affine parameter of the source is $\lambda_{s}=-|\boldsymbol{x}_{s}|/\omega_{0}$.
@@
$k^{\mu}(\lambda)=k_{(0)}^{\mu}(\lambda)+k_{(1)}^{\mu}(\lambda)+\mathcal{O}(h^{2})\;,$ (B2)
@@
$\frac{\mathrm{d}k^{\rho}}{\mathrm{d}\lambda}=-\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (B3)
We integrate the above expression from $\lambda=0$ to arbitrary $\lambda$ to obtain $k^{\mu}(\lambda)$
$k^{\mu}(\lambda)=k_{(0)}^{\mu}+\kappa^{\mu}-\int_{0}^{\lambda}\Gamma^{\rho}_{\mu\nu}k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime}\;,$ (B4)
where $\kappa^{\mu}$ is the first order perturbation of the photon four-momentum at $\lambda=0$. Integrating Eq. (B4) yields the trajectory of photon
$x^{\mu}(\lambda)=\lambda(k_{(0)}^{\mu}+\kappa^{\mu})-\int_{0}^{\lambda}\int_{0}^{\lambda^{\prime}}\Gamma^{\rho}_{\mu\nu}(\lambda^{\prime\prime})k^{\mu}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+x^{\mu}(0)\;,$ (B5)
where $x^{\mu}(0)$ is the final position of the photon, which also refers to the position of the detector. Choosing $\lambda=\lambda_{s}$ in Eq. (B5), we are able to solve $\kappa^{\mu}$
$\kappa^{\mu}=\frac{1}{\lambda_{s}}\left(x^{\mu}(\lambda_{s})-x^{\mu}(0)+\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\Gamma^{\mu}_{\gamma\nu}(\lambda^{\prime\prime})k^{\gamma}_{(0)}k^{\nu}_{(0)}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\right)-k_{(0)}^{\mu}\;.$ (B6)
To the zeroth order of perturbation, $x^{\mu}(\lambda_{s})-x^{\mu}(0)=\lambda_{s}k_{(0)}^{\mu}$. With the presence of metric perturbation, we can not assume that the detector and source are static since the geodesic equations of their spatial trajectories also contain $\mathcal{O}(h)$ perturbations. Therefore we keep $x^{\mu}(\lambda_{s})$ and $x^{\mu}(0)$ in the expression of $\kappa^{\mu}$.
In addition, the null geodesic equation constraint $g_{\mu\nu}(\lambda)k^{\mu}(\lambda)k^{\nu}(\lambda)=0$ provide
$\kappa_{0}=n_{i}\kappa_{i}-\frac{1}{2\omega_{0}}h_{\mu\nu}(0)k^{\mu}_{(0)}k^{\nu}_{(0)}\;.$ (B7)
@@
$e^{\nu}_{\alpha}=\delta^{\nu}_{\alpha}+\delta e^{\nu}_{\alpha}\;,\quad\alpha=0,1,2,3\;.$ (B8)
@@
$e^{\mu}_{0}\nabla_{\mu}e^{\nu}_{\alpha}=0\;.$ (B9)
Inserting Eq. (B8) into Eq. (B9), we find
$\partial_{0}\delta e^{\nu}_{\alpha}+\Gamma_{\alpha 0}^{\nu}=0\;.$ (B10)
@@
$g_{\mu\nu}e^{\nu}_{\alpha}e^{\mu}_{\beta}=\delta_{\alpha\beta}\;.$ (B11)
We keep only the first order terms of Eq. (B11) which gives another constraint on $\delta e^{\nu}_{\alpha}$
$h_{\mu\nu}+\eta_{\mu\gamma}\delta e^{\gamma}_{\nu}+\eta_{\nu\gamma}\delta e^{\nu}_{\mu}=0\;.$ (B12)
@@
$\delta n_{i}=-\frac{g_{\nu\mu}(0)k^{\nu}(0)e^{\mu}_{i}}{\omega_{d}}-n_{i}\;,$ (B13)
where $\omega_{d}=-g_{\nu\mu}(0)k^{\nu}$ is the observed photon frequency. Expanding Eq. (B13) to the first order, we obtain
$\delta n_{i}=-\frac{1}{\omega_{0}}\left(h_{\nu i}(0)k^{\nu}_{(0)}+\kappa_{i}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{i}-n_{i}(h_{\nu 0}(0)k^{\nu}_{(0)}+\eta_{\nu\mu}k^{\nu}_{(0)}\delta e^{\mu}_{0}+\kappa_{0})\right)\;.$ (B14)
Plugging Eq. (B7) into Eq. (B14), we find
$\delta n_{i}=-\frac{1}{\omega_{0}}P_{ij}\kappa_{j}-h_{\nu i}(0)\mathcal{N}^{\nu}-\mathcal{N}_{\mu}\delta e^{\mu}_{i}+n_{i}\left(h_{\nu 0}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}k_{d(1)}^{\mu}\right)-\frac{1}{2}n_{i}h_{\nu\mu}\mathcal{N}^{\mu}\mathcal{N}^{\nu}\;,$ (B15)
where $P_{ij}=\delta_{ij}-n_{i}n_{j}$ is the projection operator. Multiplying both sides of Eq. (B15) by $n^{i}$, we obtain
@@
$=0\;.$ (B16)
Note that in the last step we use Eq. (B12). Now $n^{i}\delta n_{i}=0$ shows that $\delta n_{i}$ is orthogonal to $n^{i}$, so we can multiply $P_{ij}$ to each term in the expression of $\delta n_{i}$, which reads
$\delta n_{i}=-P_{ij}\left(\frac{1}{\omega_{0}}\kappa_{j}+h_{\nu j}(0)\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\delta e^{\mu}_{j}\right)\;.$ (B17)